Operations Report – May 25th

Crew 195 Operations Report 25.05.18

SOL: 5

Name of person filing report: Dana Levin

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet.

Notes on non-nominal systems: n/a

Generator (hours run): Overnight approximately 8 hours, on solar until 1600, now back on generator

Solar— SOC 10%

Diesel – Over 50%

Propane – 50 PSI

ATV Fuel (Ethanol free gasoline) – 1 gallon.

Water (trailer) – 485 gallons.

Water (static) – Approximately 90

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351243

Toilet tank emptied: No
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallon

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: 1 hour
Beginning charge: 85%
Ending charge: 80 %
Currently charging: Yes

Curiosity rover used: Yes

Hours: 1 hour
Beginning charge: 99%
Ending charge: 85 %
Currently charging: Yes

Spirit rover used: no

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

GreenHab Report – May 25th

Greenhab Report

Dana Levin


Environmental control:


Shade cloth on

Average temperatures:

Low: 20C

High: 50C

Hours of supplemental light: none

Changes to crops: yup… still dead

Daily water usage for crops: n/a

Time(s) of watering for crops: n/a

Moringa research observations: n/a

Changes to research plants: n/a

Aquaponics: n/a

Narrative: we held a ceremony for the dead crops and while we were not sure what religion plants typically follow, and we were unable to ask them as they are currently dead we are confident that the great garden in the sky (or perhaps in the ground) will receive them and they shall enjoy endless sunshine filled with an abundance of rain and miracle grow. Fare well Mars crops. We think of you often…

Support/supplies needed: n/a

Sol Summary Report – May 25th

Crew 195 Daily Summary Report 25.05.2018

MDRS Daily Summary Report for sol 5

Summary Title

The final Day

Mission Status:


Sol Activity Summary:

EVA to Robert’s Rock Garden (aka the martian graveyard) was undertaken by the intrepid MDRS 195 crew. It seems that no day passes on this mission without a medical emergency and today was no different. We struggled through the crisis and the crew handled it well returning to the habitat after declaring a Delta evacuation and working through the problems before their thorough and informative debrief.

Look Ahead Plan: return to earth tomorrow

Anomalies in work:

ATV 300- continued issues with battery. Unable to hold charge.

Generator and solar continue to be manually switched

leak beneath HAB continues to occur while running pump from water to static tank.

Weather: Scattered Clouds, hot

Crew Physical Status: Doing well.


Reports to be filed:

Journalist Report

Photo report

Support Requested:


Journalist Report May 25th

Journalist Report – Earth-Date 25May2018 / Sol 5

Author: Dana Levin

This concludes the mission of crew 195. We have suffered many a crisis and the crew has weathered them bravely. We’ve been down but not out, beaten but not broken, Spurned but never scorned, cloudy but not stormy, free but not clear, and many other things. In the end, this crew has handled the rigors of our cruel mistress mars well and looks forward to our return to earth having been enriched by the experience and our lives changed permanently and forever.

EVA Report May 25th

Crew 195 EVA Report 25May2018

EVA Report


The crew set out to the nearby site they’ve dubbed the “martian graveyard” to search for signs of ancient life. Unfortunately they were yet again beset by crisis when crew members became incapacitated in the simulation. The crew expertly handled the situation as they have many times before and discussed the management and events afterwards with a fantastic debrief. This crew’s final EVA was executed with expert ability and demonstrated their command of the complex material they learned during this course.

EVA Crew Members: Scott, Madison, Rae, Christian, Dana, Mason, Barry, Ruslana,

Habitat Crew: Chris


Near Habitat

Easting: 518800

Northing: 4250900

EVA objectives

Survey or signs of life

Management of medical contingencies

Return Rovers to Hab


Medical contingencies managed

Rovers and ATVs were successfully recovered

ATVs/Rovers Used

ATVs x 3, EVs x 2 (deimos, curiosity)

Copyright © The Mars Society. All rights reserved. | Main Site